OpenStudy (anonymous):

An electric vehicle starts from rest and accelerates at a rate of 3.0 m/s2 in a straight line until it reaches a speed of 25 m/s. The vehicle then slows at a constant rate of 1.0 m/s2 until it stops. How much time elapses from start to stop?

OpenStudy (anonymous):

First Phase u=0 a=3 m/s^2 v=25 m/s Second phase u=25 m/s a=-1 m/s^2 v=o v=u+at t1=(v-u)/a t1=25/3 s t2=-25/-1 therefore total time = t1+t2 = 25+(25/3)